# The Impact of Renewable Energy Communities in the Italian Day-Ahead Electricity Market: A Scenario Analysis

### v1

@@ 4.1 Descriptive statistics of REC configurations in Italy
Table 3 displays the completeness of the available data. The information on the market zone and the type of installed technology is the most complete. Crucial data on the nominal capacity of power plants is available for 82.9% of RECs. The estimation of many parameters required data on the category of members, which is available for 74.9% of RECs. The typology of prosumer buildings is available for only 43.7% of RECs. In turn, the self-consumption level—from which the range of 45–50–55 scenarios was assumed—is available for only 22.7% of RECs. Similarly, only 3.6% of RECs plan to install BESS.
Table 3: Available data on RECs used for the study
Available Information Number of RECs
Total number of RECs 362
Information on the electricity market zone where a REC is located 362
Information on technology installed in a REC 323
Information on generator’s capacity installed in a REC 300
Information on category of members
(consumer, prosumer [public, residential, SME, NPO], producer)
271
   from which:
    Category of members and member number 219
    Only category of members without number 52
Building types where a prosumer’s facility is installed
(e.g., school, hospital, condominium, town hall, church, etc.)
158
Self-consumption level of REC prosumers 82
Battery availability and/or capacity in a REC 13
@@
In Figure 6232323Figure 6 contains information not for all 362 RECs from our database but for RECs with the available information., we observe that the most common prosumer/producer category is public prosumers—typically represented by municipal buildings. Standalone systems are reported for 61 RECs, while rooftop installations owned by SMEs and residential prosumers are reported for 42 and 36 RECs, respectively. In contrast, only 22 RECs report involving NPO prosumers. The higher number of RECs with public prosumers may be explained by the fact that municipalities are, by far, the most common promoters of RECs in Italy, for whom targeted capital and regional grants are also more accessible. In addition, public buildings consume most of their load during daylight hours, making them ideal candidates for maximizing the state incentive on shared energy. Accordingly, standalone plants also allow for the maximization of profits due to their greater average installed capacity, especially when balanced with the daytime loads of public, SME, and NPO prosumers and consumers. Although residential prosumers have not yet actively participated with their rooftop systems, many residential consumers [^24] REC consuming members are not the focus of this study. Therefore, even though data on REC consumers is available, we do not report it here. are members of RECs in Italy [^25] The elaborate discussion on equity concerns and private citizen participation in Italian RECs can be found in Koltunov (2025b)..
@@
Figure 7 shows the average capacity shares of all categories of prosumers/producers across market zones. Photovoltaic installations of all categories are present only in three zones: NORD, CSUD, and SUD. The CSUD and NORD market zones have the most proportionate prosumer/producer capacity distributions. This situation is possibly associated with the high number of RECs located in these zones (see Fig. 8). Moreover, the strong economic activity in these regions may contribute to a greater diversity of stakeholders who possess the financial, administrative, technical, legal, and social engagement resources and skills necessary to participate in RECs (Koltunov, 2025b). In contrast, a more disproportionate distribution of capacity across prosumer/producer categories - where public bodies dominate - can be observed in the market zones of Southern Italy (SUD, SARD, SICI, CALA). This differentiation into REC prosumers’ participation is also supported by the qualitative findings of Musolino et al. (2023). Importantly, the absence of installed plants for certain member categories does not automatically exclude them from REC membership. For example, all member categories may also participate in RECs as consumers without owning rooftop systems.
@@
Figure 8 demonstrates the “Average PV capacity per one Prosumer/Producer“, $P^{a.one}_{p,z}$, which is one of the two real-world input parameters for the engineering model. [^26] Table in Appendix A reports values for two constituent parameters, $P^{a.total}_{p,z}$ and $N^{r}_{z}$, that are used to derive $P^{a.one}_{p,z}$ The average capacity of a single photovoltaic plant varies across market zones. For public and SME prosumers, the average capacity in the NORD and Central zones (CNORD, CSUD) is significantly higher than in the Southern zones (CALA, SUD, SICI). [^27] Only 5 observations were used to estimate the average SME capacity in SARD. Therefore, 334.5 kW of average capacity installed by SMEs in SARD is based on a limited sample and should be interpreted with caution. Similarly, the very large capacity of public prosumers in CNORD zone, 285.5 kw, may be related to the small observational sample. Again, this pattern may be associated with easier access to financing in Northern and Central Italy. Another notable observation is that NPOs have significantly smaller installed capacities on average compared to SMEs and public prosumers. Typically, non-profit organizations have more limited access to private financing than SMEs, while capital grant subsidies are reserved exclusively for small municipalities (Koltunov, 2025b). Finally, zero values are mostly observed in the Southern zones, possibly due to the small number of REC observations.
In Figure 9, we observe the types of buildings where PV installations have been constructed by prosumer category. This information was used to identify the “most common building type.” For public prosumers, data on specific building types is unavailable for most RECs; consequently, the most common identified building type is a school. For SMEs, it is a commercial building, typically a supermarket or shopping mall. Hotels and industrial buildings are also relatively common in the SME category. The type of almost all residential buildings is unspecified (935 buildings), followed by private detached houses (14 buildings). Offices and churches are the most common building types for the NPO category. Finally, 65 standalone photovoltaic plants have been reported, while non-photovoltaic technologies have been used in a much smaller number of facilities.
@@ 4.2 Results from the synthetic approach
This subsection displays the core results of the paper, presenting main outputs at weekly, monthly, and quarterly resolutions. Results are reported exclusively for NORD and CSUD zones. [^28] Our simulation algorithm did not retrieve valid outcomes for the remaining five zones, mainly due to data discontinuities and the low diffusion of RECs in such zones.
The analysis focuses on the comparative dynamics between observed market outcomes and their synthetic counterfactuals, highlighting the extent to which REC-driven injections and self-consumption might influence zonal market equilibria in terms of traded volumes and prices.
To improve the readability of plots representing the percentage impact of RECs on equilibrium quantities in the DA market, we applied a smoothing procedure using a centered simple moving average. Specifically, for each hour of the day, we replaced the original values with the average computed over a 7-hour symmetric window, using the rollmean function from the zoo package in R. The smoothing was applied selectively, only to positive values, while non-positive entries were left unchanged. This method reduces short-term fluctuations and highlights general trends across the different scenarios and seasons. This procedure has been applied to all Figures from 14 to 23.
@@ 4.2.1 Results for main scenarios
In Figure 10, the notable difference between the magnitude of the equilibrium effects in winter (smaller spike density) and other seasons (greater spike density) can be observed. A combination of the low solar irradiance while high-self consumption in winter induces this trend. First, lower solar irradiation in winter lead to fewer quantities of energy to be injected into the grid by RECs. Second, self-consumption rate by RECs is higher during winter. Figure 11 illustrates daily self-consumption levels for all prosumer categories in different months. All categories self-consume on average more energy in January than in April, July or October. As a result, the RECs’ self-consumption effect on the demand side of the market is relatively bigger than the RECs’ injection effect on the supply side of the market in winter than in other seasons. In addition, low solar irradiance during winter aggregates the effect of offset of energy injection with energy self-consumption. In our engineering model, we added electrified heating and cooling to the load profiles of public (schools), SME (commercial), and NPO (offices) categories [^31] However, residential prosumers own just electrified cooling systems in our model and not heating systems, similar to the actual status quo in Italian households.. Therefore, in Figure 11, we observe that heating needs increase self-consumption during winter, mostly, for only three, albeit REC dominating (Figure 7), categories of prosumers.
Figure 10: Hourly impact on equilibrium quantities by RECs in the North market zone.
@@
We further investigate the latter finding, by analyzing the percentage relative difference between real and counterfactual hourly quantity profiles - stemming from the outcomes of Figure 13 -, and applying the smoothing procedure reported at the beginning of this Section. Intuitively, a positive percentage relative difference indicates that actual market quantities outweigh counterfactual synthetic volumes, while a negative percentage relative difference would signal the opposite effect.
Yet, Figure 14 shows that both the HW (sc45.HW.2027) and the BU (sc45.BU.2027) scenarios display a positive relative percentage effect of RECs on actual market quantities. Conversely, the Policy scenario (sc45.2027) falls under the zero line for several settlement periods - especially during peak hours -, thus suggesting that RECs have the potential to reduce traded market volumes in the Italian DA power market under large levels of deployment. Nevertheless, the magnitude of such effect has a strong seasonal dependency. Indeed, the largest magnitude of relative reduction in DA equilibrium quantities occurs in January, as a result of the aforementioned ”REC Winter Effect”. [^32] This entails a larger energy self-consumption by RECs relatively to their energy injection into the grid.
@@
Figure 16 shows comparable results for the HW scenario: in July, the impact of RECs on market quantities emerges earlier (4–6 a.m.), mirrors the April effect until 8 a.m., then weakens during the day, and resurfaces around 7 p.m. due to extended sunlight - similarly to the BU scenario.
As for the Policy scenario (Figure 17 ), the impact of RECs in April would oscillate around the zero-line resembling ”zig-zags” until midday, thus witnessing a balanced effect between lower market demand (due to larger self-consumption by RECs) and higher electricity supply by RECs. Instead, the impact of RECs would be more pronounced during the rest of the day. Interestingly, the effect of RECs in July is the opposite, indicating a prevailing effect of energy injected by RECs until 1 p.m., to then switch back to the dominance of the effect of reduced market demand resembling ”zig-zags”, which could be explained by the strong need of power consumption for space cooling during summer. Indeed, we notice that the green line constantly falls before dropping below zero after 2 pm. We observe that the effect of reduced market demand (explicitly evident within the Policy scenario) mimics the hourly self-consumption rates of different prosumers’ categories. For example, in July, self-consumption rates in the afternoon (1pm - 7pm) for all prosumers’ categories are higher than the morning self-consumption rates (6am - 12pm). [^33] Public prosumers raise self-consumption rates on average at 2.3%, SME prosumers at 2.7%, residential prosumers at 18.5%, and NPO prosumers at 1.1%. In contrast, in April, self-consumption rates are, on average, higher in the morning than in the afternoon. We assume that the specific composition of prosumers in RECs (Figure 7) might also affect the impact’s dynamics. However, this inquiry requires deeper investigation which is left for future research.
@@
Lastly, we break down results by showcasing weekdays and weekends effects (Figures 18-23).
@@
These latter findings carry important policy implications, particularly in light of the growing electricity demand driven by system-wide electrification and the potential role of RECs to reduce system stress. Yet, we argue that, in the scenario where RECs are widely deployed (Policy scenario), the system would recognize the role of RECs in reducing wholesale power demand through increased self-consumption. This would not only shield members of RECs from market price volatility, but also reduce the need for costly grid infrastructure investments, thereby facilitating a broader integration of RECs into the energy system, while consumers’ autonomy would be enhanced.
4.2.2 Sensitivity analysis: mixed scenarios
We hereby report a sensitivity analysis, by relaxing the assumption of the homogeneous self-consumption rate for all types of prosumers. Yet, we explore the impact of the two ”mixed” scenarios as indicated in Table 2, hence modeling heterogeneous self-consumption rates for prosumers’ categories.
@@ 4.2.2 Sensitivity analysis: mixed scenarios
Figure 24: Profile of average hourly quantities in NORD for both actual and counterfactual scenarios. Period: January 2024. Comparison between mixed scenarios: sc_\_mix1.2027 vs. sc_\_mix2.2027.
Figure 25: Hourly percentage impact on equilibrium quantities by RECs: comparison between mixed scenarios: sc_\_mix1.2027 vs sc_\_mix2.2027 for NORD. Period: January 2024.
5 Discussion, policy implications, and future research directions.
Italy allocated €5.7 billion to support REC deployment: €3.5 billion for premium tariffs (financed via a levy on electricity consumption) and €2.2 billion from the National Recovery and Resilience Fund to cover up to 40% of CAPEX in small municipalities. These figures highlight the strategic importance of assessing the effectiveness and system-wide value of RECs. Our paper delves into such issue, clarifying how trajectories of RECs deployment affect the equilibrium of the Italian DA power market. This Section presents an array of implications arising from our investigation and compare it with findings from other studies.
Firstly, our results show that the impact of RECs on power market outcomes is focused primarily in the NORD and CSUD market zones, with current negligible or absent effects in the other five physical market zones. This finding is closely related to the current distribution of REC projects, which are predominantly located in such two areas (see Fig. 8). Relatively slow and complex authorization and acceptance processes, institutional burdens, and limited technical support in other areas prevent broader deployment, confirming the need for targeted support policies to unlock REC potential across all regions. The similar finding has been reported by Zhu et al. (2025) and Musolino et al. (2023).
Secondly, the temporal dimension of REC impacts reveals strong seasonal and hourly patterns, driven by the interaction between solar production profiles and the assumed self-consumption rates. Our simulations for BU and HW scenarios indicate that certain early morning and evening hours - especially in spring and summer months - exhibit a relatively significant increase in market quantities, thereby highlighting the potential for RECs to better fulfill peak demand pressures in a decarbonizing system. Similarly, although including BESS, Riaz et al. (2019) modeled a summer week in the Australian wholesale market. They investigated the aggregate effect of a large number of prosumers on the load profile and found that increased prosumer participation with BESS flattens demand profiles, enhancing voltage stability and reducing the need for gas peaking plants. However, the exception in their study is the scenario with low demand and excess RES generation. It leads to increased aggregate demand due to battery charging, thus decreasing the stability margins of the system. Boccard and Goetz (2025) argue that excessive RES generation from REC prosumers entails increased power exchanges with the grid, although they do not specify in which direction.

### v2

@@ 4.1 Descriptive statistics of REC configurations in Italy
As we see in Table 3.1.1, the information on the market zone and the type of installed technology is the most complete (362 and 323 RECs respectively). Crucial data on the nominal capacity of power plants is available for 82.9% (300) of RECs. The estimation of many parameters required data on the share of installed PV capacity by category of prosumers, which is available for 74.9% (271) of RECs. The typology of prosumer buildings is available for 43.7% (158) of RECs. In turn, the self-consumption level—from which the range of 45–50–55 scenarios was assumed—is available for 22.7% (82) of RECs. Similarly, only 3.6% (13) of RECs plan to install BESS.
@@
In Figure 8, [^26] Figure 6 contains information not for all 362 RECs from our database but for RECs with the available information. we observe that the most common prosumer/producer category is public prosumers—typically represented by municipal buildings. Standalone systems are reported for 61 RECs, while rooftop installations owned by SMEs and residential prosumers are reported for 42 and 36 RECs, respectively. In contrast, only 22 RECs report involving NPO prosumers. The higher number of RECs with public prosumers may be explained by the fact that municipalities are, by far, the most common promoters of RECs in Italy, for whom targeted capital and regional grants are also more accessible. Moreover, public buildings consume most of their load during daylight and off-peak hours (see Fig. 4, 5), making them ideal candidates for maximizing the state incentive on shared energy (see Appendix 7.2). Accordingly, standalone plants also allow for the maximization of profits due to their greater average installed capacity, especially when balanced with the daytime loads of public, SME, and NPO prosumers and consumers. Although residential prosumers have not yet actively participated in RECs with rooftop systems, many residential consumers [^27] REC consuming members are not the focus of this study. Therefore, even though data on REC consumers is available, we do not report it here. are members of RECs in Italy [^28] The elaborate discussion on equity concerns and private citizen participation in Italian RECs can be found in Koltunov (2025b)..
@@
Figure 9 shows the average capacity shares of all categories of prosumers/producers across market zones. Photovoltaic installations of all categories are present only in three zones: NORD, CSUD, and SUD. The CSUD and NORD market zones have the most proportionate prosumer/producer capacity distributions. This situation is possibly associated with the high number of RECs located in these zones (see Fig. 10). Moreover, the strong economic activity in these regions may contribute to a greater diversity of stakeholders who possess the financial, administrative, technical, legal, and social engagement resources and skills necessary to participate in RECs (Koltunov, 2025b; Musolino et al., 2023). In contrast, a more disproportionate distribution of capacity across prosumer/producer categories—where public bodies dominate—can be observed in the market zones of Southern Italy (SUD, SARD, SICI, CALA). Importantly, the absence of installed plants for certain member categories does not automatically exclude them from REC membership. All categories of members may also participate in RECs as consumers without owning rooftop systems.
@@
Figure 10 demonstrates the “Average PV capacity per one Prosumer/Producer“, $P^{a.one}_{p,z}$, which is one of the two real-world input parameters for the engineering model. The average capacity of a single photovoltaic plant varies across market zones. For public and SME prosumers, the average capacity in the NORD and Central zones (CNORD, CSUD) is significantly higher than in the Southern zones (CALA, SUD, SICI). [^29] Only 5 observations were used to estimate the average SME capacity in SARD. Therefore, 334.5 kW of average capacity installed by SMEs in SARD is based on a limited sample and should be interpreted with caution. Similarly, the very large capacity of public prosumers in CNORD zone, 285.5 kw, may be related to the small observational sample. Another notable observation is that NPOs have significantly smaller installed capacities on average compared to SMEs and public prosumers. Typically, non-profit organizations have more limited access to private financing than SMEs, while state capital subsidies are reserved exclusively for small municipalities (Koltunov, 2025b). Finally, zero values are mostly observed in the Southern zones, possibly due to the small number of REC observations.
In Figure 11, we observe the types of buildings—where rooftop installations have been constructed—estimated for prosumer categories. This information was used to identify the “most common building type.” For public prosumers, data on specific building types is unavailable for most RECs; consequently, the most common identified building type is a school. For SMEs, it is a commercial building, typically a supermarket or shopping mall. Hotels and industrial buildings are also relatively common in the SME category. The type of almost all residential buildings is unspecified (935 buildings), followed by private detached houses (14 buildings). Offices and churches are the most common building types for the NPO category. Finally, 65 standalone photovoltaic plants have been reported, while non-photovoltaic technologies have been used in a much smaller number of facilities.
@@ 4.2 Results from the synthetic approach
This subsection displays the core results of the paper, presenting the main outputs in weekly, monthly, and quarterly resolutions. The results are reported exclusively for the NORD and CSUD zones. [^30] Our simulation algorithm did not retrieve valid results for the remaining five zones, mainly due to data discontinuities and the low diffusion of RECs in such zones, and therefore we did not include them in the paper. Nonetheless, these results can be available upon request to the corresponding author. The analysis focuses on the comparative dynamics between observed market outcomes and their synthetic counterfactuals, highlighting the extent to which REC-driven injections and self-consumption might influence zonal market equilibria in terms of traded volumes and prices. To improve the readability of the plots that represent the percentage impact of RECs on equilibrium quantities in the DA market, we applied a smoothing procedure to all Figures from 16 to 25 using a centered simple moving average. Specifically, for each hour of the day, we replaced the original values with the average computed over a 7-hour symmetric window, using the rollmean function from the zoo package in R. The smoothing was applied selectively, only to positive values, while non-positive entries were left unchanged. This method reduces short-term fluctuations and highlights general trends across the different scenarios and seasons.
@@ 4.2.1 Results for main scenarios
In Figure 12, the notable difference between the magnitude of the equilibrium effects in winter (smaller spike density) and other seasons (greater spike density) can be observed. A combination of the low solar irradiance while high-self consumption in winter induces this trend. First, lower solar irradiation in winter lead to fewer quantities of energy to be injected into the grid by RECs. Second, self-consumption rate by RECs is higher during winter. Figure 13 illustrates daily self-consumption levels for all prosumer categories in different months. All categories self-consume on average more energy in January than in April, July or October. As a result, the RECs’ self-consumption effect on the demand side of the market is relatively bigger than the RECs’ injection effect on the supply side of the market in winter than in other seasons. In our engineering model, we added electrified heating and cooling to the load profiles of public (schools), SME (commercial), and NPO (offices) categories [^34] However, residential prosumers own just electrified cooling systems in our model and not heating systems, similar to the actual status quo in Italian households.. Therefore, in Figure 13, we observe that heating needs increase self-consumption during winter, mostly, for only three, albeit REC dominating (Figure 9), categories of prosumers.
Figure 12: Hourly percentage impact on equilibrium quantities from deployment of RECs in the North market zone.
@@
We further investigate the latter finding by analyzing the percentage relative difference between average real and counterfactual hourly quantity profiles, as stemming from the outcomes displayed in Figure 15, and by applying the smoothing procedure described at the beginning of this Section. Similarly to the hourly percentage impact on equilibrium quantities defined in Eq. (22), this indicator is computed as the relative deviation between average observed and synthetic quantities over the selected time window.
Formally, for each settlement hour $h$, let $\bar{Q}_{actual,h}$ and $\bar{Q}_{synt,h}$ denote the average cleared quantities in the actual and counterfactual markets, respectively, computed over all days in the reference period. The relative difference is then defined as:
$\text{RelDiff}_{h}=\frac{\bar{Q}_{actual,h}-\bar{Q}_{synt,h}}{\bar{Q}_{synt,h}}\times 100.$ (23)
By construction, a positive relative difference indicates that, on average, market equilibrium volumes are higher in the presence of RECs compared to the no-REC baseline, while a negative value signals a contraction in traded quantities, typically associated with a prevailing effect of REC self-consumption. This metric enables a transparent assessment of both the magnitude and the temporal (intra-day and seasonal) heterogeneity of REC-driven deviations from counterfactual market equilibria.
Yet, Figure 16 shows that both the HW (sc45.HW.2027) and the BU (sc45.BU.2027) scenarios display a positive relative percentage effect of RECs on actual market quantities. Conversely, the Policy scenario (sc45.2027) falls under the zero line for several settlement periods, thus suggesting that RECs have the potential to reduce traded market volumes in the Italian DA power market under large levels of deployment. Nevertheless, the magnitude of such effect has a strong seasonal dependency. Indeed, the largest magnitude of relative reduction in DA equilibrium quantities occurs in January during off-peak hours (10 a.m.–14 p.m.), as a result of the aforementioned ”REC winter effect”. These reductions during off-peak hours are related to loads from institutional REC members (public, SME, NPO).
@@
As for the Policy scenario (Figure 19 ), the impact of RECs in April would oscillate around the zero-line resembling ”zig-zags” until midday, thus witnessing a balanced effect between lower market demand (due to larger self-consumption by RECs) and higher electricity supply by RECs. Instead, the impact of RECs would be more pronounced during the rest of the day. Interestingly, the effect of RECs in July is the opposite, indicating a prevailing effect of energy injected by RECs until 1 p.m., to then switch back to the dominance of the effect of reduced market demand resembling ”zig-zags”, which could be explained by the strong need of power consumption for space cooling during summer. Indeed, we notice that the green line constantly falls before dropping below zero after 2 pm. We observe that the effect of reduced market demand (explicitly evident within the Policy scenario) mimics the hourly self-consumption rates of different prosumers’ categories. For example, in July, self-consumption rates in the afternoon (1pm - 7pm) for all prosumers’ categories are higher than the morning self-consumption rates (6am - 12pm). [^35] Public prosumers raise self-consumption rates on average at 2.3%, SME prosumers at 2.7%, residential prosumers at 18.5%, and NPO prosumers at 1.1%. In contrast, in April, self-consumption rates are, on average, higher in the morning than in the afternoon. We think that the specific composition of prosumers in RECs (Figure 9) can also affect the impact’s dynamics.
@@
4.2.2 Sensitivity analysis: mixed scenario
@@
In the scenario where RECs are widely deployed (Policy scenario), the system would recognize the role of RECs in reducing wholesale power demand through increased self-consumption. This would not only shield members of RECs from market price volatility, but also reduce the need for costly grid infrastructure investments while consumers’ autonomy would be enhanced.
Figure 26: Profile of average hourly quantities in NORD for both actual and counterfactual scenarios. Period: January 2027. Comparison between mixed scenario and policy scenario: sc_\_mix1.2027 vs. sc55.2027.
Figure 27: Hourly percentage impact on equilibrium quantities by RECs: comparison between mixed scenario and 3d policy scenario: sc_\_mix1.2027 vs sc55.2027 for NORD. Period: January 2027.
5 Discussion
First, our results show that the impact of RECs on power market is focused primarily in the NORD and CSUD market zones, with current negligible or absent effects in the other five physical market zones. This finding is closely related to the current distribution of REC projects, which are predominantly located in such two areas (see Fig. 10). Relatively slow and complex authorization and acceptance processes, institutional burdens, and limited technical support in other areas prevent broader deployment, confirming the need for targeted support policies to unlock REC potential across all regions. The similar finding has been reported by Zhu et al. (2025) and Musolino et al. (2023).
Second, the temporal dimension of REC impacts reveals strong seasonal and hourly patterns, driven by the interaction between solar production profiles and the assumed self-consumption rates. Our simulations for BU and HW scenarios indicate that certain early morning and evening hours - especially in spring and summer months - exhibit a relatively significant increase in market quantities, thereby highlighting the potential for RECs to better fulfill peak demand pressures in a decarbonizing system. Similarly, although including BESS, Riaz et al. (2019) modeled a summer week in the Australian wholesale market. They investigated the aggregate effect of a large number of prosumers on the load profile and found that increased prosumer participation with BESS flattens demand profiles, enhancing voltage stability and reducing the need for gas peaking plants. However, the exception in their study is the scenario with low demand and excess RES generation. It leads to increased aggregate demand due to battery charging, thus decreasing the stability margins of the system. Boccard and Goetz (2025) argue that excessive RES generation from REC prosumers entails increased power exchanges with the grid, although they do not specify in which direction.